OpenStudy (anonymous):

The hypotenuse of a right triangle is 50 mm long. One leg of the right triangle is 30 mm long. What is the length of the other leg? 20mm 30mm 40mm 60mm

OpenStudy (anonymous):

40

OpenStudy (anonymous):

40

OpenStudy (lalaly):

40

OpenStudy (angela210793):

x^2=50^2-30^2=2500-900=1600 x=sqr1600=40